OpenStudy (anonymous):

a snowball melts so tha the surface area decreases at a rate of 1cm^2/min. the surface area of the sphere with radius of a(r)=pi(r)^2. find an equation relating the surface area of the snowball to its diamater d

OpenStudy (anonymous):

Notice how the Surface Area of a sphere is:\[S.A=4 \pi r^2\] This can be re-written as:\[S.A=(2r)^2 \pi\]We also know that the diameter is twice the radius or \(d=2r\). So now we re-write the equation as:\[S.A=\pi (d)^2\] @djt0206